# Altermagnetic phase transition in a Lieb metal

### v1

@@ A Static four-point TUFRG
We use the FRG in its static four-point approximation. This means that we disregard all higher (six, eight, …) point interaction vertices as well as self-energies in the expansion of the vertex generating functional [30]. The resulting diagrammatic structure of the FRG flow equations is visualized in Fig. S1. In addition to the trunctation of the number of fermionic fields participating in possible interactions, we disregard the frequency dependence of all vertices. This approximation has proven useful when characterizing ordering propensities of weakly to intermediately interacting electron systems [29]. As the four-point vertices are still complicated objects, they are routinely compressed in numerical simulations: The primary (bosonic) momentum is resolved in momentum space, while the secondary (ferminoic) momenta are treated in a truncated real-space basis. This coins the “truncated unity” approximation of FRG, i.e., TUFRG [40, 41, 42, 43].
@@ C Mean-field altermagnetic phase transition from FRG
S4 Stability analysis of the altermagnetic phase
@@ A Interaction parameters
[…]
@@
Figure S3: (a) Ferromagnetic (FM) phase emerging at small $t^{\prime}$. As the $A$ site participates in the magnetization process, a slight tendency towards ferrimagnetism can be observed. (b) $XY$-spin density wave emerging at the interface of FM and AM. At the phase transition, all linear combinations of the twofold degenerate order parameter are allowed solutions of the linearized gap equation.
@@ B Kinetic parameters
Figure S4 displays different values of intrinsic detuning $\mu_{A}\neq\mu_{B,C}=0.0$ and NNN hopping $t^{\prime}$ at pristine filling. The sublattice polarized VHS at the $X$/$Y$ point persists for arbitrary values of $\mu_{A}$ and arbitrary but finite values of $t^{\prime}$ as it corresponds to a topologically protected band touching point. This property is reflected in the resilience of the AM phase upon variation of $t^{\prime}\geq 0.5$ (see Fig. S2b) and $\mu_{A}=-2.0\dots 3.0$.
@@
For the given band structure, itinerant phases are generally suppressed for small amounts of hole doping as the DOS rapidly declines (see Fig. 1b). For small amounts of electron doping, the $B$/$C$ polarized band is still present at the Fermi surface stabilizing the AM phase with tendencies to form the particle-hole condensate at slightly incommensurate momenta. As the logarithmically diverging DOS and perfect nesting scenario is absent at the Fermi level, intermediate interaction values are necessary to obtain the altermagnetic phase in our FRG calculations (e.g., $U_{i}=1.4\dots\geq 2.6$ for $\mu=0.02$).
@@ C Dominance of the altermagnetic order parameter
Figure S5: Bandstructure and eigenvalues of the bare susceptibility at lower VHS (left) and middle VHS (right) for the Hamiltonian in Eq. 1 of the main text for $t=-1$, $t^{\prime}=0.5t$ and $\mu_{A}=0$. We evaluated the bare susceptibility via Eq. S20 on a $\bm{k}$-mesh of $2000\times 2000$ and set $\beta=100/t$. The insets shows the eigenvector associated with the leading eigenvalues at distinct $\bm{Q}$ values in the whole BZ. We note that ferromagnetic fluctuations are suppressed on the bare level already when doping to the center VHS (left column).
S5 Symmetry classification of the $d$-wave altermagnetic order
Non-relativistic collinear magnetism is generally classified by spin groups [20, 18] described as the direct product: $r_{s}\times R_{s}$. Here, $r_{s}$ refers to the spin only group whose implication on the band structure of collinear magnets is $\epsilon_{\sigma}(k_{x},k_{y})=\epsilon_{\sigma}(-k_{x},-k_{y})$ regardless of whether the real-space inversion symmetry is present. Here, $r_{s}=[C_{\infty}||E]+[\bar{C}_{2}C_{\infty}||E]$ where $C_{\infty}$ is any rotation around the common spin axis and $\bar{C}_{2}$ is the two-fold rotation around the axis perpendicular to the spins combined with spin space inversion. In the notation $[R_{i}||R_{j}]$, $R_{i}(R_{j})$ incorporates all the spin (crystallographic) space operations.
@@
The observed phase in this study belongs to the third category. Specifically, the non-trivial spin group is given by $[C_{2}||A]$ where $A$ includes three symmetry elements $\mathcal{M}_{xy}$ (mirror plane perpendicular to $xy$ plane and passing through $x=y$), $\mathcal{M}_{x\bar{y}}$ (mirror plane perpendicular to $xy$ plane and passing through $x=-y$) and $C_{4}$ ($C_{4z}$ rotational axis) that interchanges B and C sublattices:
::: $\epsilon_{\sigma}(k_{x},k_{y})\rightarrow\epsilon_{-\sigma}(k_{y},k_{x})\,,$ (S22)
@@
S6 Size of the non-relativistic spin splitting
@@
with $\Delta_{M}$ ($-\Delta_{M}$) the spin polarization on the $B$ ($C$) sublattice. The corresponding eigenvalues for spin-$\downarrow$ are obtained by flipping the sign: $\Delta_{M}\rightarrow-\Delta_{M}$.
@@
S7 Realization in an optical lattice
@@
Figure S7: Possibility of realization Lieb system in an optical lattice simulator. (a) Optical lattice potential given by Eq. S28 with $(v_{1},v_{2},r)=(4.0,4.0,0.7)$, (b) Experimental setup. (c) top panels: optical lattice band structure (single particle) for $(v_{1},v_{2})$ given by $(4.1,4.1)$, $(4.25,4.25)$ and $(4.1,4.1)$ respectively and $r=0.7$. We consider the energy offset as $E_{0}=26E_{R}$. (c) bottom panels: Tight binding band structure with $(t^{\prime},\mu_{A})$ given by $(0.2,-0.3)$,$(0.2,0.0)$ and $(0.2,0.3)$ in Eq. 1 where $\mu_{A}$ and $t^{\prime}$ are scaled with $t$.

### v2

@@ A Static four-point TUFRG
We use the FRG in its static four-point approximation. This means that we disregard all higher (six, eight, …) point interaction vertices as well as self-energies in the expansion of the vertex generating functional [50]. The resulting diagrammatic structure of the FRG flow equations is visualized in Fig. S1. In addition to the trunctation of the number of fermionic fields participating in possible interactions, we disregard the frequency dependence of all vertices. This approximation has proven useful when characterizing ordering propensities of weakly to intermediately interacting electron systems [64]. As the four-point vertices are still complicated objects, they are routinely compressed in numerical simulations: The primary (bosonic) momentum is resolved in momentum space, while the secondary (ferminoic) momenta are treated in a truncated real-space basis resulting in three distinct diagrammatic channels (particle-particle ($P$), crossed particle-hole ($C$), and direct particle-hole ($D$)). This coins the “truncated unity” approximation of FRG, i.e., TUFRG [29, 42, 59, 3].
The renormalization group flow is terminated upon encountering a divergent vertex element at some critical cutoff $\Lambda_{c}$ indicating an instability of the Fermi liquid towards a symmetry broken phase. Since all quantum fluctuations above $\Lambda_{C}$ are included in the effective vertex, the residual effective interaction can be analyzed within a simple MF theory to reveal the expected symmetry breaking [62, 56]. In order to carry out the Hubbard-Stratonovich transformation of the effective 4-point vertex, a fermionic bilinear has to be chosen as static MF. There are three possible choices, corresponding to superconducting ($\langle c^{\dagger}_{\bm{k}^{\prime},\uparrow}\,c^{\dagger}_{\bm{q}-\bm{k}^{\prime},\downarrow}\rangle$), charge ($\langle c^{\dagger}_{\bm{k}^{\prime},\uparrow}\,c^{\phantom{\dagger}}_{\bm{q}+\bm{k}^{\prime},\uparrow}+c^{\dagger}_{\bm{k}^{\prime},\downarrow}\,c^{\phantom{\dagger}}_{\bm{q}+\bm{k}^{\prime},\downarrow}\rangle$) or magnetic ($\langle c^{\dagger}_{\bm{k}^{\prime},\uparrow}\,c^{\phantom{\dagger}}_{\bm{q}+\bm{k}^{\prime},\uparrow}-c^{\dagger}_{\bm{k}^{\prime},\downarrow}\,c^{\phantom{\dagger}}_{\bm{q}+\bm{k}^{\prime},\downarrow}\rangle$) instability. The MF interaction matrices $V$ for the respective orders can be directly obtained from the interaction in the diagrammatic channels depicted in Fig. S1 as $V^{\text{SC}}=V^{P}$, $V^{\text{charge}}=2V^{D}-V^{C}$, and $V^{\text{magnetic}}=V^{C}$.
For the physical interaction channels a MF decoupling is performed by defining the non-vanishing expectation value i.e. in the magnetic channel as
$\Delta_{\bm{q},\,1\,4}(\bm{k})\,=\,\sum_{\bm{k}^{\prime},3\,2}\,V^{\text{magnetic}}_{1\,4\,3\,2}(\bm{q},\bm{k},\bm{k}^{\prime})\langle c^{\dagger}_{\bm{k}^{\prime},\uparrow,3}\,c^{\phantom{\dagger}}_{\bm{k}^{\prime}+\bm{q},\uparrow,2}-\,c^{\dagger}_{\bm{k}^{\prime},\downarrow,3}\,c^{\phantom{\dagger}}_{\bm{k}^{\prime}+\bm{q},\downarrow,2}\rangle,$ (S9)
where all quantum numbers other than spin are combined in roman numeral multi-indices. Since the FGR flow is only valid until the phase transition, it is meaningful to apply the MF theory directly at the phase boundary where $\Delta\rightarrow 0$. In this limit the self-consistent gap equation ca be linearized and reduces to a simple eigenvalue equation, that ca be evaluated independently for every transfer momentum $\bm{q}$
$\lambda_{\bm{q}}\,\Delta_{\bm{q},\zeta}\,=\,\sum_{\zeta^{\prime}}V_{\zeta\,\zeta^{\prime}}(\bm{q})\,\Delta_{\bm{q},\zeta^{\prime}}\,$, (S10)
where $\lambda_{\mathbf{q}}$ acts as a proxy for the transition temperature for the respective order. The physically realized instability is then indicated by the eigenvector $\Delta$ associated with the largest eigenvalue $\lambda$ over all channels and wave vectors $\mathbf{q}$. Here, the vertices are already represented in the matrix basis, absorbing the reordering complexity within the index $\zeta\leftrightarrow(1\,4\,\bm{k})$ and $\zeta^{\prime}\leftrightarrow(3\,2\,\bm{k}^{\prime})$.
@@ C Mean-field altermagnetic phase transition from FRG
S4 Long range magnetic order in 2D
It is well known, that two dimensional spin systems lack true long range order due to the Mermin-Wagner-Hohenberg (MWH) theorem at any finite temperature [49, 27]. While it predicts a finite correlation length for the spin-spin expectation value, this length scale can be exponentially large. Combined with the omnipresent coupling of any real two dimensional system with the three dimensional environment in physical conditions renders the implications of the MWH theorem irrelevant for practical use cases.
In the present theoretical model, that considers a real two dimensional material, we would nevertheless expect the MWH theorem to hold. MF theory does not capture quantum fluctuation responsible for the destruction of long range coherence and is hence not expected to comply with the MWH theorem. The FRG on the other hand takes into account fluctuations in an unbiased manner and still predicts a divergence of a collective spin excitation mode at finite temperature (or likewise critical cutoff) at odds with MWH. This is a well known flaw of the 1-loop approximation of the FRG flow equations visualized in Fig. S1 required to conveniently solve the infinite hierarchy of coupled differential equations [13]. It has been shown, that the multi-loop extension to the FRG is able to reproduce exact solutions in accordance with MWH [39, 26]. However, these extensions are not applicable to multi-orbital systems due to their immense computational costs. Still, the FRG in its 1-loop approximation is able to qualitatively predict the most dominant collective excitations, that will form quasi-long range order, that can be identified with a true long range coherence for all practical purposes even in the strict 2D limit.
S5 Stability analysis of the altermagnetic phase
@@ A Interaction parameters
We demonstrate that the proposed AM state remains robust across the entire weak to intermediate coupling regime of onsite interactions $U_{A}=U_{B,C}$. Figure S2 a shows the critical cutoff scale $\Lambda_{c}$ for different values of $U$. The cutoff scale, being a handle on the critical temperature, increases rapidly with increased coupling strength. Furthermore, we analyze how different onsite interactions on the two types of sublattices affect the stability of the state, i.e. $U_{A}\neq U_{B,C}$. For fixed $U_{B,C}=1$ the AM state persists for $U_{A}=0\,\dots\,2.5$.
@@
Figure S3: (a) Ferromagnetic (FM) phase emerging at small $t^{\prime}$. As the $A$ site participates in the magnetization process, a slight tendency towards ferrimagnetism can be observed. (b) XYXY-spin density wave emerging at the interface of FM and AM. At the phase transition, all linear combinations of the twofold degenerate order parameter are allowed solutions of the linearized gap equation.
@@ B Kinetic parameters
Figure S4 displays different values of intrinsic detuning $\mu_{A}\neq\mu_{B,C}=0.0$ and NNN hopping $t^{\prime}$ at pristine filling. The sublattice polarized VHS at the XX/$Y$ point persists for arbitrary values of $\mu_{A}$ and arbitrary but finite values of $t^{\prime}$ as it corresponds to a topologically protected band touching point. This property is reflected in the resilience of the AM phase upon variation of $t^{\prime}\geq 0.5$ (see Fig. S2b) and $\mu_{A}=-2.0\dots 3.0$.
@@
For the given band structure, itinerant phases are generally suppressed for small amounts of hole doping as the DOS rapidly declines (see Fig. 1b). For small amounts of electron doping, the BB/$C$ polarized band is still present at the Fermi surface stabilizing the AM phase with tendencies to form the particle-hole condensate at slightly incommensurate momenta. As the logarithmically diverging DOS and perfect nesting scenario is absent at the Fermi level, intermediate interaction values are necessary to obtain the altermagnetic phase in our FRG calculations (e.g., $U_{i}=1.4\dots\geq 2.6$ for $\mu=0.02$).
@@ C Dominance of the altermagnetic order parameter
The dominant peak in the bare susceptibility not only allows to identify the mot prominent ph fluctuations in the system, but directly transfers to the transition temperature in a general MF study of the system. To see this, we can perform a generic mean field decomposition of the free energy in Eq. S12 and expand it up to second order in the magnetic order parameters $\vec{\Delta}$ to obtain the Ginzburg-Landau (GL) functional
$F_{\text{GL}}=\sum_{\mathbf{Q}}\Big[\frac{1}{U}-\chi^{0}(\mathbf{Q})\Big]\vec{\Delta}(\mathbf{Q})^{2}+\mathcal{O}(\vec{\Delta}^{4})\ ,$ (S24)
where we sum over different potentially degenerate wave vectors $\mathbf{Q}$ and dropped the orbital and spin indices of $\vec{\Delta}$ and $\chi^{0}$ brevity. This approximation is valid directly at the phase transition where $\vec{\Delta}\rightarrow 0$ and requires more and more terms as one enters the ordered phase. The expected transition temperature $T_{c}$ is given by the temperature for which the determinant of the second order coefficient matrix. In the case of a single ordering vector, $T_{c}$ is determined by the condition $1/U=\chi^{0}(\mathbf{Q})$, where the bare susceptibility is temperature dependent as evident from Eq. S22 and grows as $T\rightarrow 0$. Since we expect no change in the hierarchy of eigenstates as T is lowered, the eigenvalues of $\chi^{0}(\mathbf{Q})$ in Fig. S5 give a direct estimate for the free energy gain in the symmetry broken phase for $T<T_{c}$. Consequently, we expect already at the MF level the AM state to provide the predominant instability for the flat band at the Fermi level. It is also worth noting, that the first terms in the GL expansion, that allows a coupling of the different magnetic states present in Fig. S2 is of fourth order due to the different irreducible representations of the AM ($B_{1}$), FM ($A_{1}$) and $XY$-SDW ($\mathbf{Q}=X/Y$) order parameter
$F_{\text{GL}}^{4}=\vec{\Delta}_{\text{AM}}^{2}\vec{\Delta}_{i}(\mathbf{Q}_{i})^{2}+(\vec{\Delta}_{\text{AM}}\cdot\vec{\Delta}_{i}(\mathbf{Q}_{i}))^{2}\,$ (S25)
where $\vec{\Delta}_{i}(\mathbf{Q}_{i})\in\{\vec{\Delta}_{\text{FM}}(\mathbf{Q}=0),\vec{\Delta}_{\text{SDW}}(\mathbf{Q}=X/Y)\}$. Therefore, we do not expect a sizable admixture of other magnetic states in the AM phase. This validates the single component contrained MF calculation of Section C.
Figure S5: Bandstructure and eigenvalues of the bare susceptibility at lower VHS (left) and middle VHS (right) for the Hamiltonian in Eq. 1 of the main text for t=1t=-1, $t^{\prime}=0.5t$ and $\mu_{A}=0$. We evaluated the bare susceptibility via Eq. S22 on a $\bm{k}$-mesh of $2000\times 2000$ and set $\beta=100/t$. The insets shows the eigenvector associated with the leading eigenvalues at distinct $\bm{Q}$ values in the whole BZ. We note that ferromagnetic fluctuations are suppressed on the bare level already when doping to the center VHS (left column).
S6 Symmetry classification of the $d$-wave altermagnetic order
Non-relativistic collinear magnetism is generally classified by spin groups [72, 85] described as the direct product: $r_{s}\times R_{s}$. Here, $r_{s}$ refers to the spin only group whose implication on the band structure of collinear magnets is $\epsilon_{\sigma}(k_{x},k_{y})=\epsilon_{\sigma}(-k_{x},-k_{y})$ regardless of whether the real-space inversion symmetry is present. Here, $r_{s}=[C_{\infty}||E]+[\bar{C}_{2}C_{\infty}||E]$ where $C_{\infty}$ is any rotation around the common spin axis and $\bar{C}_{2}$ is the two-fold rotation around the axis perpendicular to the spins combined with spin space inversion. In the notation $R_{i}||R_{j}]$, $R_{i}(R_{j})$ incorporates all the spin (crystallographic) space operations.
@@
The observed phase in this study belongs to the third category. Specifically, the non-trivial spin group is given by $C_{2}||A]$ where $A$ includes three symmetry elements $\mathcal{M}_{xy}$ (mirror plane perpendicular to $xy$ plane and passing through $x=y$), $\mathcal{M}_{x\bar{y}}$ (mirror plane perpendicular to $xy$ plane and passing through x=yx=-y) and $C_{4}$ ($C_{4z}$ rotational axis) that interchanges B and C sublattices:
$:$ $\epsilon_{\sigma}(k_{x},k_{y})\rightarrow\epsilon_{-\sigma}(k_{y},k_{x})\,$, (S26)
@@
S7 Size of the non-relativistic spin splitting
@@
with $\Delta_{M}$ ($\Delta_{M}$) the spin polarization on the $B$ ($C$) sublattice. The corresponding eigenvalues for spin-$\downarrow$ are obtained by flipping the sign: $\Delta_{M}\rightarrow-\Delta_{M}$.
@@
S8 Realization in an optical lattice
@@
Figure S7: Possibility of realization Lieb system in an optical lattice simulator. (a) Optical lattice potential given by Eq. S32 with $(v_{1},v_{2},r)=(4.0,4.0,0.7)$, (b) Experimental setup. (c) top panels: optical lattice band structure (single particle) for $(v_{1},v_{2})$ given by $(4.1,4.1)$, $(4.25,4.25)$ and $(4.1,4.1)$ respectively and $r=0.7$. We consider the energy offset as $E_{0}=26E_{R}$. (c) bottom panels: Tight binding band structure with $(t^{\prime},\mu_{A})$ given by (0.2,0.3)(0.2,-0.3),$(0.2,0.0)$ and $(0.2,0.3)$ in Eq. 1 where $\mu_{A}$ and $t^{\prime}$ are scaled with $t$.